# Dynamic Programming: From Local Optimality to Global Optimality

### v2

@@ Appendix A Algorithm
Input: Model parameters $(\beta,\gamma)$, distributions $(\psi,\varphi)$, episodes $K$, batch size $N$
Output: Policy network $\hat{\sigma}$, Value network $\hat{v}_{\sigma}$
1
2Initialize policy network $\hat{\sigma}(\cdot;\theta)$ and value network $\hat{v}_{\sigma}(\cdot;\lambda)$;
3 Initialize policy optimizer with learning rate $\alpha_{\theta}$ and value optimizer with learning rate $\alpha_{\lambda}$;
4 Initialize initial states in batch $w\in\mathbbm{R}^{N}$;
5
6for episode $k=1$ to $K$ do
       $\tilde{w}\leftarrow(w-w_{\min})/(w_{\max}-w_{\min})$ ;
        // Normalize states
@@ 4 Initialize initial states in batch $w\in\mathbbm{R}^{N}$;
      $c\leftarrow\hat{\sigma}(\tilde{w};\theta)$ ;
        // Current policy
       $v\leftarrow\hat{v}_{\sigma}(\tilde{w};\lambda)$ ;
        // Value estimate
       $r\leftarrow u(c)$ ;
        // Current reward
8
      $\eta\sim\psi$, $y\sim\varphi$ ;
        // Sample shocks
       $w^{\prime}\leftarrow\eta(w-c)+y$ ;
        // Next state
9       $\tilde{w}^{\prime}\leftarrow(w^{\prime}-w_{\min})/(w_{\max}-w_{\min})$ ;
@@
      $v^{\prime}\leftarrow\hat{v}_{\sigma}(\tilde{w}^{\prime};\lambda)$ ;
        // Next state value
       $v_{\text{target}}\leftarrow r+\beta v^{\prime}$ ;
        // Bellman target
11
      $L\leftarrow\frac{1}{N}\sum_{i=1}^{N}(v_{i}-v_{\text{target},i})^{2}$ ;
        // Bellman error
       $\lambda\leftarrow\lambda-\alpha_{\lambda}\nabla_{\lambda}L$ ;
        // Value update
12
      $J\leftarrow\frac{1}{N}\sum_{i=1}^{N}(r_{i}+\beta\hat{v}_{\sigma}(\tilde{w}^{\prime}_{i};\lambda))$ ;
        // Expected return
       $\theta\leftarrow\theta+\alpha_{\theta}\nabla_{\theta}J$ ;
        // Policy improvement
13
      $w\leftarrow w^{\prime}$ ;
        // State evolution
14
@@
16return $\hat{\sigma},\hat{v}_{\sigma}$;
Algorithm 1 Deep Policy Value Iteration for Optimal Savings
Appendix B Irreducible Optimal Savings MDP at $w=1$
Figure 6. $\hat{v}_{\sigma}$ and $\hat{\sigma}$ with $w=1$ against the OPI solutions.
Figure 7. Expected returns (13) and value loss (14) over training episode for the irreducible optimal saving model at $w=1$.

### v3

@@ Appendix A Algorithm
This section provides a detailed description of the algorithm used to train the policy network described in Section 5.1. In the state transition step, the algorithm uses the wealth bounds $(w_{\min},w_{\max})$ to ensure that the wealth process remains within these bounds to match the OPI solution, which operates on a wealth grid. In practice, we can replace it with $w_{t+1}\leftarrow\eta_{t}(w_{t}-c)+y_{t}$.
We implement the policy network and optimizer using PyTorch. The training is conducted on an NVIDIA RTX 5070 Ti GPU with CUDA 12.8. Our implementation uses Python 3.13.2 and leverages PyTorch’s built-in automatic differentiation (Paszke et al., 2017) to compute gradients of the loss function efficiently.
Input: Discount factor $\beta$, distributions $(\psi,\varphi)$, episodes $K$, rollout steps $T$, batch size $N$, learning rate $\alpha$, wealth bounds $(w_{\min},w_{\max})$
1 Initialize policy network $\hat{\sigma}$ and optimizer
2 for episode $k=1$ to $K$ do
3       Initialize states $w\leftarrow\bar{w}\in\mathbbm{R}^{N}$
4       Pre-sample shocks vectors $\eta_{t}\sim\psi$, $y_{t}\sim\varphi$ for all $t=0,1,\ldots,T-1$
5       for $t=0$ to $T-1$ do
6             $c\leftarrow\hat{\sigma}(w_{t};\theta)$
             $w_{t+1}\leftarrow\min\{\max\{\eta_{t}(w_{t}-c)+y_{t},w_{\min}\},w_{\max}\}$
              // Transitions
@@ 6             $c\leftarrow\hat{\sigma}(w_{t};\theta)$
8       end for
9      $\theta\leftarrow\theta-\alpha\nabla_{\theta}L(\theta)=\theta+\alpha\nabla_{\theta}\frac{1}{N}\sum_{i=1}^{N}\sum_{t=0}^{T-1}\beta^{t}u(\hat{\sigma}(w_{i,t};\theta))$
@@
return $\hat{\sigma}(\cdot;\theta)$
Algorithm 1 Deterministic Policy Gradient
Appendix B Irreducible Optimal Savings MDP with $\bar{w}=50$
Figure 6. $\hat{v}_{\sigma}$ and $\hat{\sigma}$ with $\bar{w}=50$ against the OPI solutions.
Figure 7. $\hat{v}_{\hat{\sigma}}(\bar{w})$ over training episode for the irreducible optimal saving model at $\bar{w}=50$.
Appendix C Reducible Optimal Savings MDP with $\bar{w}=50$
In this appendix, we discuss the result of reducible optimal saving MDPs with initial wealth $\bar{w}=50$. As shown in Figure 9, the policy network $\hat{\sigma}$ is able to approximate the optimal policy $\sigma^{*}$ in the reachable subset of the state space. This provides strong empirical evidence for Lemma 4.1.
To emphasize this point, Figure 10 illustrates one sample of the wealth dynamics under the policy network $\hat{\sigma}$ at epoch 750 starting from $\bar{w}=1$ and $\bar{w}=50$. The wealth process initialized at $\bar{w}=50$ explores a substantially broader range of wealth levels compared to the process starting from $\bar{w}=1$. This wider exploration of the state space directly contributes to the improved performance observed in Figure 9.
We note that the two sequences converge to similar patterns after a brief initial period because we used identical random seeds when generating shocks $\eta_{t}$ and $y_{t}$ when training all models. This technical choice minimizes the influence of factors other than the initial wealth level on the comparisons of model performance. The similarity in the wealth trajectories indicates that the policy network gives similar consumption decisions at lower wealth levels.
Nevertheless, the policy still exhibits limitations: since consumption must be strictly positive (i.e., $\hat{\sigma}(w)>0$ for all $w\in\mathsf{W}$), the algorithm struggles to accurately approximate the optimal policy for very high wealth values which are not reachable under $\bar{w}=50$ given the bounded shocks. This discrepancy paired with bounded uniform shocks causes the $\hat{\sigma}$-value function to systematically underestimate the true value function in the upper tail of the wealth state in the left panel of Figure 9.
Figure 8. $\hat{v}_{\sigma}$ and $\hat{\sigma}$ with $\bar{w}=50$ vs OPI solutions
Figure 9. $\hat{v}_{\hat{\sigma}}(\bar{w})$ over training episode for the reducible optimal saving model at $\bar{w}=50$.
These empirical findings emphasize the importance of Lemma 4.1, which formally establishes that when the irreducibility condition fails, the policy network achieves near-optimal performance in reachable subsets of the state space starting from $\bar{w}=50$.
Figure 10. Comparison of wealth trajectories for $\bar{w}=1$ and $\bar{w}=50$ in a typical rollout.